OpenStudy (me123456):

Please help with step by step answers to solving the problem! The sum of the first n terms of an arithmetic sequence with first term a1 and a common difference d is Sn. Determine the sum if each term is increased by 5.

OpenStudy (anonymous):

sn=[n(a1+an)/2] an=a1+(n-1)d Find your a1; d is given.

OpenStudy (jit4won):

If you increase each term by 5 it means you are adding 5 to every single term

OpenStudy (jit4won):

Total number of terms = n so you have added 5 ,n times i.e You added 5n sum is sn +5n

OpenStudy (me123456):

I'm still confused on how writing out the formula to figure out the sum.

OpenStudy (anonymous):

I think you are missing some information. All you are given is d, but not a1. You cant really find an, or sn without more info

OpenStudy (me123456):

It's like proving the theory of arithmetic sequence and finding its sum.

OpenStudy (jit4won):

n/2(2a1+(n-1)d) + 5n

OpenStudy (me123456):

how did you get the 2?

OpenStudy (jit4won):

sn +5n =n/2(2a1+(n-1)d) + 5n because sn =n/2(2a1+(n-1)d)

OpenStudy (jit4won):

You wanna me to prove sn= n/2(2a1+(n-1)d) ?

OpenStudy (me123456):

No, I get it sorta now. If I have any other questions I'll ask.

OpenStudy (me123456):

Isn't the formula for the sum of arithmetic sequence supposed to be n(a1+an)/2?

OpenStudy (jit4won):

yeah But l= a1+(n-1)d

OpenStudy (anonymous):

formulas are in the first post

OpenStudy (me123456):

so d = 5?

OpenStudy (anonymous):

yes

OpenStudy (jit4won):

Nope

OpenStudy (jit4won):

d= d

OpenStudy (jit4won):

even if each term is increased by 5 ,the common difference won't change you should change you first term .put a1 +5 at the place of a1

OpenStudy (jit4won):

you will get n/2(2a1+(n-1)d) + 5n

OpenStudy (me123456):

Can you write it out step by step and show me?

OpenStudy (jit4won):

Sure my if my AP is a1 , a1+ d , a1+ 2d ,............................... a1+(n-1)d sn=n/2(2a1+(n-1)d) If I add 5 to every every number AP becomes a1+5 , a1+d+5, a1+2d+5 ,................................ a1+(n-1)d +5 instead of a1 we'll put a1+5 because our 1st term is now a1+5 check d hast change yet sn =n/2(2(a1+5)+(n-1)d) =n/2(2a1+(n-1)d) + 5n

OpenStudy (jit4won):

Sure my if my AP is a1 , a1+ d , a1+ 2d ,............................... a1+(n-1)d sn=n/2(2a1+(n-1)d) If I add 5 to every number, AP becomes a1+5 , a1+d+5, a1+2d+5 ,................................ a1+(n-1)d +5 instead of a1 we'll put a1+5 because our 1st term is now a1+5 check d hasn't change yet sn =n/2(2(a1+5)+(n-1)d) =n/2(2a1+(n-1)d) + 5n

OpenStudy (me123456):

Thanks.

OpenStudy (jit4won):

You're welcome .... I'm always ready to help those who want my help